# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ 4.2 The Semantics of $S^{\star}$ in the SCM with Interventions Setting
The results in Section 4.1 are for the problem nonparametric invariance pursuit itself. In a population-level view, it shows that if $S^{\star}$ satisfies 4.2 and 4.5, or in other (informal) words, $S^{\star}$ is the “maximum invariant set” in that it satisfies
@@
$m^{(e,S^{\star})}\equiv\bar{m}^{(S^{\star})}~{}(\text{invariant})\qquad\text{and}\qquad\forall S\subseteq[d],~{}m^{(e,S)}\equiv\bar{m}^{(S)}\Longrightarrow\bar{m}^{(S\cup S^{\star})}=\bar{m}^{(S^{\star})}~{}(\text{maximum})$
simultaneously, the sample-efficiency estimation can be attained by our proposed FAIR-NN estimator. It is natural to ask
@@
This section is devoted to offering a clean yet general answer to the problem under the SCM with intervention (on $X$) setting. The short answer is: Yes, and it is a type of data-driven causality.
To start with, we first introduce the concept of the structural causal model (Glymour et al.,, 2016).
@@
As in Peters et al., (2016), we consider the following data-generating process in $|\mathcal{E}|$ environments. For each $e\in\mathcal{E}$, the process governing $p=d+1$ random variables $Z^{(e)}=(Z_{1}^{(e)},\ldots,Z_{d+1}^{(e)})=(X_{1}^{(e)},\ldots,X_{d}^{(e)},Y^{(e)})$ is derived from an SCM $M^{(e)}(\mathcal{S}^{(e)},\nu)$, whose induced graph $G(M^{(e)})$ is acyclic, and structural assignments are as follows,
@@
Here the distribution of exogenous variables $(U_{1},\ldots,U_{d+1})$, the cause-effect relationship $G$, and the structural assignment $f_{d+1}$ are invariant across $e\in\mathcal{E}$, while the structural assignments for $X$ may vary among $e\in\mathcal{E}$. We use superscript $(e)$ to highlight this heterogeneity. This heterogeneity may arise from performing arbitrary intervention on the variables $X$. We use $Z_{\mathtt{pa}(j)}$ to emphasize that $Y$ can be the direct cause of some variables in the covariate vector. See an example in Fig. 1 (a).
To present the result, we consider an augmented SCM that incorporates the environment label $e$ as a variable $E$. We consider the case where =$\mathcal{E}=\{0,\ldots,|\mathcal{E}|-1\}$. We let 00 be the observational environment, and the rest are the interventional environments where some unknown, arbitrary interventions are applied to the variables in some given set (but is agnostic to our estimation method) $I\subseteq[d]$. The shared cause-effect relationships in all the environments are encoded by $G$, or $\{\mathtt{pa}(j)\}_{j=1}^{d}$, based on which we define $\mathtt{ch}(k):=\{j:k\in\mathtt{pa}(j)\}$ as the set of children of variable $k$ and $\mathtt{at}(k)$ as the set of all the ancestors of the variable $Z_{k}$, defined recursively as $\mathtt{at}(k)=\mathtt{pa}(k)\cup\cup_{j\in\mathtt{pa}(k)}\mathtt{at}(j)$ in the topological order of $G$.
We can define a unified SCM $\widetilde{M}=(\widetilde{\mathcal{S}},\widetilde{\nu})$ on $d+2$ variables $Z=(Z_{1},\ldots,Z_{d},Z_{d+1},Z_{d+2})=(X_{1},\ldots,X_{d},Y,E)$ encoding all the information of $|\mathcal{E}|$ models $M^{(0)}(\mathcal{S}^{(0)},\nu),\ldots,M^{(|\mathcal{E}|-1)}(\mathcal{S}^{(|\mathcal{E}|-1)},\nu)$ defined in (4.4). Here $\widetilde{\nu}(du_{1},\ldots,du_{d+2})=\nu(du_{1},\ldots,du_{d+1})\nu_{b}(du_{d+2})$ with $\nu_{b}\sim\mathrm{Uniform}(\mathcal{E})$, and the structural assignments $\widetilde{\mathcal{S}}=\{\widetilde{f}_{1},\ldots,\widetilde{f}_{d+2}\}$ are defined as follows:
$\begin{split}E&\leftarrow\widetilde{f}_{d+2}(U_{d+2}):=U_{d+2}\\ X_{j}&\leftarrow\begin{cases}\widetilde{f}_{j}(X_{\mathtt{pa}(j)},U_{j}):=f^{(0)}_{j}(X_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in[d]\setminus I\\ \widetilde{f}_{j}(X_{\mathtt{pa}(j)},E,U_{j}):=f^{(E)}_{j}(X_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in I\end{cases}\\ Y&\leftarrow\widetilde{f}_{d+1}(X_{\mathtt{pa}}(d+1),U_{d+1}):=f_{d+1}(X_{\mathtt{pa(d+1)}},U_{d+1}).\end{split}$ (4.5)
It should be noted that throughout this section, the direct cause map $\mathtt{pa}:[d+1]\to[d+1]$ matches the causal relationship $G$ instead of $\widetilde{G}=G(\widetilde{M})$. See a graphical illustration of the above construction in Fig. 1.
@@
Figure 1: (a) is an illustration of the two-environment model, the SCMs in the two environments share the same associated graph, that is also plotted in (b): $M^{(0)}$ is an observational environment, and $M^{(1)}$ is an interventional environment where some unknown intervention is applied to $(X_{4},X_{6},X_{7})$, $M^{(0)}$ and $M^{(1)}$ are defined as (4.4). (b) visualizes $\widetilde{G}$, the associated graph of $\widetilde{M}$ constructed based on $(M^{(0)},M^{(1)})$ and (4.5).
We summarize the above construction as a condition.
Condition 4.7 (SCM with Interventions on $X$).
@@
This section aims to characterize (1) what $S^{\star}$ would satisfy (1.1) and 4.5 given a fixed intervention set $I$, and (2) how large $I$ should be to recover the $Y$’s direct causes under arbitrary types of interventions. To this end, we need the following condition ruling out some degenerated cases.
Condition 4.8 (Nondegenerate Interventions).
@@
$\forall~{}\text{Disjoint}~{}A,B,C\subseteq[d+2],\qquad Z_{A}\perp\!\!\!\perp Z_{B}|Z_{C}~{}~{}\overset{(a)}{\Longrightarrow}~{}~{}Z_{A}\perp\!\!\!\perp_{\widetilde{G}}Z_{B}|Z_{C},$
@@
The condition (2), faithfulness on the graph $\widetilde{G}$ constraining that the graph $\widetilde{G}$ truly depicts all the conditional independence relationships, is widely used in the causal discovery literature. Condition (1) is further imposed since we only leverage the information of conditional expectations instead of conditional distributions. We impose 4.8 such that the dependence on $E$ in conditional expectation of $Y$ given $X_{S}$ with any $S\subseteq[d]$ can be represented by the graph $\widetilde{G}$ itself. The imposed 4.8 rules out the possibility of some degenerated cases; see the justifications for 4.8 and some degenerated examples in Section A.5. It should be noted that our general results in Theorem 4.3 and Proposition 4.5 apply to arbitrary forms of interventions under 4.8, which is a mild condition as violation of faithfulness in 4.8 occurs with probability 00 under some suitable measure on the model (Spirtes et al.,, 2000).
Theorem 4.3 (General Identification under SCM with Interventions on $X$).
Under 4.7, there exists some set $S_{\star}$, defined as
@@
with $A(I)=\{j:j\in\mathtt{ch}(d+1),\mathtt{at}(j)\cap\mathtt{ch}(d+1)\cap I=\emptyset\}$ such that $m^{(e,S_{\star})}\equiv\bar{m}^{(S_{\star})}:=m_{\star}$. Suppose further 4.8 holds, then 4.5 holds with $S^{\star}=S_{\star}$ and $m^{\star}=m_{\star}$.
@@
Theorem 4.3 exactly characterizes what $S^{\star}$ in our nonparametric invariance pursuit in (1.1) is in the setting of SCM with interventions on $X$ for general and nondegenerate intervention. Notably, it doesn’t require intervention to be “sufficient”. Firstly, such a $S^{\star}$ is well-defined in that there exists one maximum set $S_{\star}$ satisfying the invariance condition (1.1) and heterogeneity condition 4.5 simultaneously. Secondly, in the SCM setting, such a $S^{\star}=S_{\star}$ can be represented in a simple way in (4.6), which lies in between the Markov blanket of the variable $Y$ and the set of $Y$’s direct causes. Note $A(I)$ can be interpreted as the “unaffected” children of $Y$ surviving from the interventions $I$. Then $S^{\star}$ is the union of (1) parents of $Y$, (2) unaffected children of $Y$; and (3) parents of these unaffected children. The size of that set $S^{\star}$ will keep decreasing when $I$ enlarges, and it will recover the set of direct causes of $Y$ when $I$ covers “root children set” $I^{\star}$ as stated in the following Proposition 4.5.
Unfortunately, $S_{\star}\supsetneq\mathtt{pa}(d+1)$ when $I^{\star}\not\subseteq I$ in general. However, $S_{\star}$ in this setup can be interpreted as the “contemporary direct causes” or “pragmatic direct causes” of $Y$ based on the observed environments. If the future interventions are made within the set $I$, then $S_{\star}$ can be regarded as the direct causes since the conditional expectation of $Y$ given $X_{S_{\star}}$ will remain invariant in a new environment $t$. Moreover, one can deploy such a predictor in unseen environments because it depicts the most predictive one among all the associations in environment 00 that remains in environment $t$. This can be formally stated in Proposition 4.4.
Proposition 4.4 (Robust Transfer Learning).
Under 4.7, for a new environment $t$ with SCM $M^{(t)}=\{\mathcal{S}^{(t)},\nu\}$ satisfying $f_{j}^{(t)}\equiv f_{j}^{(0)}$ for any $j\in[d+1]\setminus I$, i.e., only $X_{I}$ is intervened, we have $\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]\equiv\mathbb{E}[Y^{(0)}|X_{S_{\star}}^{(0)}]$ with $S_{\star}$ in (4.6). If 4.8 holds and $M^{(t)}$ satisfies a condition akin to 4.8 (see the details in Section A.6), then $S_{\star}$ is the maximum set whose conditional expectation is transferable in that for any $S\subseteq[d]$ such that $\mathbb{E}[Y^{(t)}|X_{S_{\star}\cup S}^{(t)}]\neq\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]$, one has $\mathbb{E}[Y^{(t)}|X_{S}^{(t)}]\neq\mathbb{E}[Y^{(0)}|X_{S}^{(0)}]$.
Proposition 4.5 (Direct Cause Recovery).
@@
(Sufficiency) Under 4.7, define $I^{\star}=\{j:j\in\mathtt{ch}(d+1),\mathtt{pa}(j)\cap\mathtt{ch}(d+1)=\emptyset\}$. If 4.8 holds and $I\supseteq I^{\star}$, then 4.5 holds with $S^{\star}=\mathtt{pa}(d+1)$.
@@
(Necessity) Moreover, if $\bar{m}^{(S^{\star}\cup S)}\neq m^{\star}$ for any $S$ with $\{j:d+1\in\mathtt{pa}(j)\}\cap S\neq\emptyset$, i.e., $Y$ does not have degenerated children, then 4.5 holds only if $I\supseteq I^{\star}$.
@@
We refer to $I^{\star}$ as the minimal intervention set because it is the exact minimal set of variables that should be intervened on for identification in general, nondegenerated cases. The set $I^{\star}$ is determined by the cause-effect relationship $G$. In particular, $I^{\star}$ is $\{6,7\}$ for the example in Fig. 1. Notably, $X_{8}$ does not require intervention, as $X_{7}$, one of its direct causes, is included in $I^{\star}$.
5 Sketch of the Generic Result and Its Applications

### v2

@@ 3 Nonparametric Invariance Pursuit under SCMs
The results in Section 2 are for the problem nonparametric invariance pursuit itself. In a population-level view, it pursues “maximum invariant set” $S^{\star}$ satisfying
@@
$m^{(e,S^{\star})}\equiv\bar{m}^{(S^{\star})}~{}(\text{invariant})~{}~{}~{}\text{and}~{}~{}~{}\forall S\subseteq[d],~{}m^{(e,S)}\equiv\bar{m}^{(S)}\Longrightarrow\bar{m}^{(S\cup S^{\star})}=\bar{m}^{(S^{\star})}~{}(\text{maximum}).$ (3.1)
Section 2 shows the FAIR-NN estimator can estimate such a $S^{\star}$ ($m^{\star}$) efficiently. It is natural to ask
@@
We offer a clean yet general answer to the question under the SCM with arbitrary interventions (on $X$) setting. The short answer is: Yes, and it can be interpreted as the “pragmatic direct causes”.
3.1 Structural Causal Model with Interventions on Covariates
We first introduce the concept of the structural causal model (Glymour et al.,, 2016). See Fig. 1 for examples of SCM. It says that each variable in the directed graph is a function of its parents (if any) and an independent innovation or noise.
@@ 3.1 Structural Causal Model with Interventions on Covariates
As in Peters et al., (2016), we consider the following data-generating process in $|\mathcal{E}|$ environments. For each $e\in\mathcal{E}$, the process governing $p=d+1$ random variables $Z^{(e)}=(Z_{1}^{(e)},\ldots,Z_{d+1}^{(e)})=(X_{1}^{(e)},\ldots,X_{d}^{(e)},Y^{(e)})$ is derived from an SCM $M^{(e)}(\mathcal{S}^{(e)},\nu)$, whose induced graph $G(M^{(e)})$ is acyclic, and assignments as
@@
Here the distribution of exogenous variables $(U_{1},\ldots,U_{d+1})$, the cause-effect relationship graph $G$, and the structural assignment $f_{d+1}$ are invariant across $e\in\mathcal{E}$, while the structural assignments for $X$ may vary among $e\in\mathcal{E}$. We use superscript $(e)$ to highlight this heterogeneity. This heterogeneity may arise from performing arbitrary interventions on the variables $X$. We use $Z_{\mathtt{pa}(j)}$ to emphasize that $Y$ can be the direct cause of some variables in the covariate vector. See an example in Fig. 1 (a).
To present the result, we consider an augmented SCM that incorporates the environment label $e$ as a variable $E$. We consider the case where =$\mathcal{E}=\{0,\ldots,|\mathcal{E}|-1\}$. We let 00 be the observational environment, and the rest are the interventional environments where some unknown, arbitrary interventions are applied to the variables in some given set $I\subseteq[d]$ defined as $I:=\{j:\exists e\in\mathcal{E}~{}\text{s.t.}~{}f_{j}^{(e)}\neq f_{j}^{(0)}\}$. The interventions can be arbitrary: it can be a “hard” do-intervention via set $X_{j}$ being $v_{j}$, or soft intervention that slightly perturbs the association, e.g., replace $X_{j}\leftarrow 2X_{k}+U_{j}$ by $X_{j}\leftarrow 1.5X_{k}+U_{j}$. The shared cause-effect relationships in all the environments are encoded by $G$, or $\{\mathtt{pa}(j)\}_{j=1}^{d+1}$.
The following SCM $\widetilde{M}=(\widetilde{\mathcal{S}},\widetilde{\nu})$ on $d+2$ variables $Z=(Z_{1},\ldots,Z_{d},Z_{d+1},Z_{d+2})=(X_{1},\ldots,X_{d},Y,E)$ encodes all the information of $|\mathcal{E}|$ models $\{M^{(e)}(\mathcal{S}^{(e)},\nu)\}_{e\in\mathcal{E}}$ in (3.2). Denote $\nu_{b}\sim\mathrm{Uniform}(\mathcal{E})$. Here $\widetilde{\nu}(du_{1},\ldots,du_{d+2})=\nu(du_{1},\ldots,du_{d+1})\nu_{b}(du_{d+2})$, and the assignments $\widetilde{\mathcal{S}}=\{\widetilde{f}_{1},\ldots,\widetilde{f}_{d+2}\}$ are defined as
$\begin{split}E&\leftarrow\widetilde{f}_{d+2}(U_{d+2}):=U_{d+2}\\ X_{j}&\leftarrow\begin{cases}\widetilde{f}_{j}(Z_{\mathtt{pa}(j)},U_{j}):=f^{(0)}_{j}(Z_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in[d]\setminus I\\ \widetilde{f}_{j}(Z_{\mathtt{pa}(j)},E,U_{j}):=f^{(E)}_{j}(Z_{\mathtt{pa}(j)},U_{j})&\qquad\forall j\in I\end{cases}\\ Y&\leftarrow\widetilde{f}_{d+1}(X_{\mathtt{pa}}(d+1),U_{d+1}):=f_{d+1}(X_{\mathtt{pa(d+1)}},U_{d+1}),\end{split}$ (3.3)
where $I$ is the set of all intervention variables in $\mathcal{E}$. It should be noted that throughout this section, the direct cause map $\mathtt{pa}:[d+1]\to[d+1]$ matches the causal relationship $G$ instead of $\widetilde{G}=G(\widetilde{M})$. See a graphical illustration of the construction in Fig. 1 (b).
We summarize the above construction as a condition.
@@
[…]
Figure 1: (a) is an illustration of the two-environment model, the SCMs in the two environments share the same associated graph: $M^{(0)}$ is an observational environment, and $M^{(1)}$ is an intervention environment where some unknown intervention is applied to $(X_{4},X_{6},X_{7})$, where $M^{(0)}$ and $M^{(1)}$ are defined as (3.2). (b) visualizes $\widetilde{G}$, the associated graph of $\widetilde{M}$ constructed based on $(M^{(0)},M^{(1)})$ and (3.3), which is another plot of the environments in (a). (c) An illustration of Theorem 2 by showing how $S_{\star}$ therein will change as we see more and more environments: the arrow from $E$ to $X_{j}$ with color $e$ means $X_{j}$ is intervened in $e\in\{{\color[rgb]{0.01953125,0.203125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.01953125,0.203125,0.546875}1},{\color[rgb]{0.8984375,0.66015625,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.8984375,0.66015625,0.29296875}2},{\color[rgb]{0.41796875,0.703125,0.57421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.41796875,0.703125,0.57421875}3},{\color[rgb]{0.68359375,0.09765625,0.03125}\definecolor[named]{pgfstrokecolor}{rgb}{0.68359375,0.09765625,0.03125}4}\}$. For example, $0\leftrightarrow{\color[rgb]{0.41796875,0.703125,0.57421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.41796875,0.703125,0.57421875}3}$ means with interventions in environments 1, 2, and 3, the invariant variable set is ${\color[rgb]{0.41796875,0.703125,0.57421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.41796875,0.703125,0.57421875}\{1,2,3,7\}}$. Although $X_{7}$ and is reverse causal and hence related to $Y$, we do not know this based only on the given environments.
Condition 4 (SCM with Interventions on $X$).
@@
3.2 Maximum Invariant Set as the Pragmatic Direct Causes
We characterize what $S^{\star}$ would satisfy (3.1) given a fixed intervention set $I$, and how large $I$ should be to recover the $Y$’s direct causes under arbitrary types of interventions. We define $\mathtt{ch}(k):=\{j:k\in\mathtt{pa}(j)\}$ as the set of children of variable $k$ and $\mathtt{at}(k)$ as the set of all the ancestors of the variable $Z_{k}$, defined recursively as $\mathtt{at}(k)=\mathtt{pa}(k)\cup\cup_{j\in\mathtt{pa}(k)}\mathtt{at}(j)$ in the topological order of $G$. The following condition rules out some degenerated cases.
Condition 5 (Nondegenerate Interventions).
@@ 3.2 Maximum Invariant Set as the Pragmatic Direct Causes
$\forall~{}\text{Disjoint}~{}A,B,C\subseteq[d+2],\qquad Z_{A}\perp\!\!\!\perp Z_{B}|Z_{C}~{}~{}{\Longrightarrow}~{}~{}Z_{A}\perp\!\!\!\perp_{\widetilde{G}}Z_{B}|Z_{C},$
@@
The condition (b), faithfulness on the graph $\widetilde{G}$ constraining that the graph $\widetilde{G}$ truly depicts all the conditional independence relationships, is widely used in the causal discovery literature. Condition (a) is further imposed since we only leverage the information of conditional expectations instead of conditional distributions. We impose 5 such that the dependence on $E$ in conditional expectation of $Y$ given $X_{S}$ with any $S\subseteq[d]$ can be represented by the graph $\widetilde{G}$ itself. The imposed 5 rules out the possibility of some degenerated cases; see the justifications for 5 and some degenerated examples in Section A.5. It should be noted that our general results in Theorem 2 and Proposition 1 apply to arbitrary forms of interventions under 5, which is a mild condition as the violation of faithfulness in 5 occurs with probability zero under some suitable measure on the model (Spirtes et al.,, 2000).
Theorem 2 (General Identification under SCM with Interventions on $X$).
Under 4, for
@@
with $A(I)=\{j:j\in\mathtt{ch}(d+1),\mathtt{at}(j)\cap\mathtt{ch}(d+1)\cap I=\emptyset\}$, we have the invariance $m^{(e,S_{\star})}\equiv\bar{m}^{(S_{\star})}:=m_{\star}$. Suppose further 5 holds, then 2 holds with $S^{\star}=S_{\star}$ and $m^{\star}=m_{\star}$.
@@
Theorem 2 exactly characterizes what $S^{\star}$ is in our nonparametric invariant pursuit under the SCM with interventions on $X$ – it doesn’t require intervention to be “sufficient”. Firstly, such a $S^{\star}$ is well-defined in that there exists one maximum set $S_{\star}$ satisfying the invariant condition (1.1) and heterogeneity condition 2 simultaneously. Secondly, in the SCM setting, such a $S^{\star}=S_{\star}$ can be represented in a simple way in (3.4), which lies in between the Markov blanket of the variable $Y$ and the set of $Y$’s direct causes. Note that $A(I)$ can be interpreted as the “unaffected” children of $Y$ from the interventions $I$. Theorem 2 states explicitly that the pursued set of invariant variables $S^{\star}$ is the union of (1) parents of $Y$, (2) unaffected children of $Y$; and (3) parents of these unaffected children. The size of that set $S^{\star}$ will keep decreasing when $I$ enlarges. It will finally recover the set of direct causes of $Y$ when $I$ includes “root children set” $I^{\star}$ as stated in the following Proposition 1. See an illustration in Fig. 1 (c).
Proposition 1 (Direct Cause Recovery).
@@
(Sufficiency) Under 4, define $I^{\star}=\{j:j\in\mathtt{ch}(d+1),\mathtt{at}(j)\cap\mathtt{ch}(d+1)=\emptyset\}$. If 5 holds and $I\supseteq I^{\star}$, then 2 holds with $S^{\star}=\mathtt{pa}(d+1)$.
@@
(Necessity) Moreover, if $\bar{m}^{(S^{\star}\cup S)}\neq m^{\star}$ for any $S$ with $\mathtt{ch}(d+1)\cap S\neq\emptyset$, i.e., $Y$ does not have degenerated children, then 2 holds only if $I\supseteq I^{\star}$.
@@
We refer to $I^{\star}$ as the minimal intervention set because it is the exact minimal set of variables that should be intervened on for exact direct cause recovery in general, nondegenerated cases. The set $I^{\star}$ is determined by the cause-effect relationship graph $G$. In particular, $I^{\star}$ is $\{6,7\}$ for the example in Fig. 1. Notably, $X_{8}$ does not require intervention, as $X_{7}$, one of its ancestors, is included in $I^{\star}$.
Unfortunately, $S_{\star}\supsetneq\mathtt{pa}(d+1)$ when $I^{\star}\not\subseteq I$ in general. This is due to a lack of evidence in environments to falsify that some variables in $S^{\star}$ are not direct causes. Nevertheless, $S_{\star}$ in this setup can still be interpreted as the “contemporary direct causes” or “pragmatic direct causes” of $Y$ based on the observed environments. If the future interventions are made within the set $I$, then $S_{\star}$ can be regarded as the direct causes since the conditional expectation of $Y$ given $X_{S_{\star}}$ will remain invariant in a new environment $t$. Moreover, one can deploy such a predictor in unseen environments because it depicts the most predictive one among all the associations in environment 00 that remains in environment $t$. This can be formally stated in Proposition 2.
Proposition 2 (Robust Transfer Learning).
Under 4, for a new environment $t$ with SCM $M^{(t)}=\{\mathcal{S}^{(t)},\nu\}$ satisfying $f_{j}^{(t)}\equiv f_{j}^{(0)}$ for any $j\in[d+1]\setminus I$, i.e., only $X_{I}$ is intervened, we have $\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]\equiv\mathbb{E}[Y^{(0)}|X_{S_{\star}}^{(0)}]$ with $S_{\star}$ in (3.4). If 5 holds and $M^{(t)}$ satisfies a condition akin to 5 (see Section A.6), then $S_{\star}$ is the maximum set whose conditional expectation is transferable in that for any $S\subseteq[d]$ such that $\mathbb{E}[Y^{(t)}|X_{S_{\star}\cup S}^{(t)}]\neq\mathbb{E}[Y^{(t)}|X_{S_{\star}}^{(t)}]$, one has $\mathbb{E}[Y^{(t)}|X_{S}^{(t)}]\neq\mathbb{E}[Y^{(0)}|X_{S}^{(0)}]$.
4 A Unified Framework
The proposed FAIR-NN least squares is a special instance of our generic FAIR estimation framework, which homogenizes different risk loss and prediction models. Moreover, our framework also allows the user to incorporate additional structural knowledge into estimation such that identification is sometimes viable when $|\mathcal{E}|=1$. The invariance pursuit problem, the estimation method, and the non-asymptotic results will be presented in a unified manner in this section.
4.1 General Invariance Pursuit from Heterogeneous Environments
In this section, we formalize the problem of invariance pursuit using data from multiple environments, which admits the canonical nonparametric invariance pursuit in Section 1.1 as a special case.
$\forall e\in\mathcal{E}\qquad\mathbb{E}\left[\left(Y^{(e)}-g^{\star}(X_{S^{\star}}^{(e)})\right)f(X^{(e)}_{S^{\star}})\right]=0\qquad\forall f\in[\Theta_{f}]_{S^{\star}},$ (4.1)
where $S^{\star}$ is the unknown set of true important variables. We refer to the above problem as invariance pursuit or causal pursuit exchangeably, as no evidence against casualty with the available experiments.
The problem of estimating $g^{\star}$ in (4.1) is a generalized version of the canonical nonparametric invariance pursuit with $g^{\star}=m^{\star}$ in (1.1) and $\Theta_{f}=\Theta_{g}=\Theta$. It depicts a general form and unifies several problems of interest in predecessors. For example, when $\Theta_{g}$ and $\Theta_{f}$ are all linear function classes, it reduces to the linear invariance pursuit problem, i.e., estimating $g^{\star}(x)=(\beta^{\star})^{\top}x=(\beta^{\star}_{S^{\star}})^{\top}x_{S^{\star}}$ with $\beta^{\star}\in\mathbb{R}^{d}$ satisfying $\mathrm{supp}(\beta^{\star})=S^{\star}$ in the multi-environment linear regression (Fan et al.,, 2023) with linear invariance structure
Another example is the augmented linear invariance pursuit where $\Theta_{g}$ is linear and $\Theta_{f}=\{f(x)=\sum_{j=1}^{d}\beta_{0,j}x_{j}+\beta_{1,j}\phi(x_{j})\}$ with some transform function $\phi:\mathbb{R}\to\mathbb{R}$. This can further generalize this to multiple transformed testing functions such as $\phi_{1}(x_{j})=x_{j}^{2}$ and $\phi_{2}(x_{j})=|x_{j}|$ but we keep one here for simplicity. The augmented linear invariance structure that realizes (4.1) in this case is
$\mathbb{E}\left[\left(Y^{(e)}-(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}\right)X_{j}^{(e)}\right]=\mathbb{E}\left[\left(Y^{(e)}-(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}\right)\phi(X_{j}^{(e)})\right]=0\qquad\forall e\in\mathcal{E},j\in S^{\star}$ (4.3)
It coincides with the problem considered by Fan & Liao, (2014) when $|\mathcal{E}|=1$ and our method reduces to the FGMM method therein. The augmented linear invariance pursuit leverages further a part of the structural knowledge that $\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]=(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}$, which is much weaker than the assumption $\mathbb{E}[Y^{(e)}|X^{(e)}]=(\beta^{\star}_{S^{\star}})^{\top}X_{S^{\star}}^{(e)}$ in the sparse linear regression. Identification is possible in this case even when $|\mathcal{E}|=1$. This is important for most biological medical studies, where data are usually collected in similar settings. In this case, the FAIR penalty eliminates endogenous spurious variables, making traditional variable selection methods applicable.
Remark 7.
We point out here that there are two kinds of spurious variables. One is endogenous spurious variables such as $X_{2}=$ background color, and the other is exogenous spurious variables such as $X_{3}=$ the time the photo was taken or the types of camera used. The former is harmful, and the latter is nearly harmless in statistical prediction, transfer learning, and even statistical attribution or causality, thinking of $X_{3}$ as a weak causal variable. The introduction of our FAIR method is to surely screen the endogenous spurious variables (Fan & Lv,, 2008). Exogenous spurious variables can be reduced by using commonly used statistical variable selection methods.
Similar to the discussion in Section 1.1, the main challenge here is the curse of endogeneity. To address this issue, we will harness the insight that the distributions of $(X,Y)$ across diverse environments capture the invariance structure (4.1). The central idea of this paper is to exploit both the heterogeneity among different environments, i.e., the shifts in population distributions $\mu^{(e)}$, in conjunction with the above invariance structure (4.1) to pinpoint the invariant regression function $g^{\star}$.
It should be noted that both $g^{\star}$ and $S^{\star}$ are determined by $(\Theta_{g},\Theta_{f})$ and $\mathcal{E}$ through the structure (4.1). It is required that $\partial\Theta_{g}=\{g-g^{\prime}:g,g^{\prime}\in\Theta_{g}\}\subseteq\Theta_{f}$. In the case of $\Theta_{f}=\partial\Theta_{g}$, one uses only heterogeneity among different environments, or the “invariance principle”, to identify the invariant regression function $g^{\star}$, as in (4.2). Heterogeneous environments are essential in this case. By choosing substantially large $\Theta_{f}\supsetneq\partial\Theta_{g}$, one further injects the strong structural assumption that the invariant regression function lies in the class $\Theta_{g}$ rather than $\Theta_{f}\setminus\Theta_{g}$ as in (4.3). In this case, one leverages both heterogeneity among environments, i.e., the “invariance principle”, and the mentioned prior structure knowledge, i.e., the “asymmetry principle”, to jointly identify $g^{\star}$. Only one environment may be enough for identifying $g^{\star}$ when the intersection of both principles gives sufficient conditions.
4.2 General FAIR Estimation Framework
$\frac{\partial\ell(y,v)}{\partial v}=(v-y)\psi(v)\qquad\text{and}\qquad\frac{\partial^{2}\ell(y,v)}{\partial v^{2}}>0,$ (4.4)
which is slightly more general than the quasi-likelihood in the generalized linear model (Nelder & Wedderburn,, 1972). The constraints in (4.4) ensure that the conditional expectation aligns with the unique global minima and can be satisfied by various risk losses. Two leading examples are the least square loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ with $\psi(v)=1$ for regression, and the cross-entropy loss $\ell(y,v)=-\log(1-v)-y\log\{v/(1-v)\}$ with $\psi(v)=1/\{v(1-v)\}$ for classification.
where $\mathcal{G}\subseteq\Theta_{g}$ and $\mathcal{F}\subseteq\Theta_{f}$ are function classes that approximates $\Theta_{g}$ and $\Theta_{f}$, respectively. Here $\widehat{\mathsf{R}}(g)$ is the pooled sample mean of the user-specified loss across all the environments $\mathcal{E}$:
$\widehat{\mathsf{R}}(g)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\widehat{\mathbb{E}}\left[\ell(Y^{(e)},g(X^{(e)}))\right]=\frac{1}{|\mathcal{E}|\cdot n}\sum_{e\in\mathcal{E},i\in[n]}\ell(Y^{(e)}_{i},g(X^{(e)}_{i})),$ (4.6)
$\gamma$ is the hyper-parameter to be determined, and $\widehat{\mathsf{J}}(g,f^{\mathcal{E}})$ is defined the same as (2.2).
4.3 Sketch of the Generic Result and Its Applications